# From Dashcam Videos to Driving Simulations: Stress Testing Automated Vehicles against Rare Events

### v1

@@ Introduction
In this paper, we propose a novel framework that automates the conversion of real-world driving videos into detailed simulation scenarios. Our approach leverages prompt-engineered Video Language Models(VLMs) to transform dashcam videos into SCENIC scripts, enabling the automatic generation of realistic simulations in CARLA. In addition, rather than aiming solely to create an exact digital copy of the dashcam video, our framework prioritizes capturing the most essential driving behaviors while maintaining flexibility in parameters such as weather and road conditions. This flexibility supports search-based testing, allowing for diverse variations of the original scenario to be explored systematically. Furthermore, we introduce a similarity metric that iteratively refines the generated simulations by comparing key driving features between the real and simulated videos. The contributions of this work are fourfold: (1) an automated video-to-simulation pipeline that removes the need for manual scenario construction, (2) a similarity metric that bridge the gap between real and simulated scenarios, (3) an iterative feedback loop for scenario refinement using neural network-based feedback from VLM to capture core driving behaviors, and (4) a significant improvement in time efficiency, reducing scenario generation from hours to minutes while maintaining high fidelity to the original events.
Related Work
@@
In contrast, our framework integrates real-world driving data, such as dashcam crash videos, to generate simulation scenarios that not only capture critical driving behaviors but also provide flexibility for environmental parameters like road type and weather. Through prompt engineering, we tailor VLMs to create SCENIC scripts for scenario generation, bridging the gap between exact scenario replication and the need for diversity in simulation environments. This approach ensures realistic and behaviorally accurate test cases while supporting search-based testing for robust ADS validation.
Real-to-Sim Scenario Generation Framework
Given a real-life vehicle crash video (e.g., a dash camera recording), our objective is to generate corresponding simulation scenarios that accurately capture the core driving behaviors. This approach enables robust and safe testing of Automated Driving Systems (ADS) by leveraging real-life close-call situations within a controlled simulation environment. Our framework consists of 4 components: (1) conversion of real-world video into SCENIC scripts, (2) generation of simulation videos from SCENIC scripts, (3) similarity analysis between the real and simulated videos, and (4) iterative refinement to ensure the simulated video’s consistency with the original scenario.
Video-to-Text Generation
@@
First, we convert the input video into a descriptive script. This is accomplished using a prompt-engineered version of the pre-trained GPT-4o model.
@@
Although the initial prompt engineering process was conducted using simulation videos paired with their SCENIC scripts, this approach is extendable to real-world driving videos because the underlying visual and descriptive patterns (e.g., road layouts, traffic behaviors, and environmental factors) are consistent across both domains, allowing the model to effectively generalize its learned capabilities and accurately capture real-world scenarios.
Text-to-Video Generation
@@
Similarity Check
Next, we perform a similarity check on the simulated scenario $V_{sim}$ and the original $V_{real}$. Ideally, they should closely match, allowing us to seamlessly replace the ego vehicle with any ADS (e.g. Baidu’s Apollo planner (Fan et al. 2018) and controller) for testing. However, due to the complexity of real-world scenarios, even with prompt engineering, discrepancies often arise, where the generated simulation may miss key features or introduce extra ones.
To address this, we introduce a similarity metric to ensure that the generated video captures the most important features from the original video. We predefine a set of crucial feature categories, such as the most critical and frequently encountered driving behaviors and environmental conditions, to examine the original crash video $V_{real}$ with the generated simulation $V_{sim}$. Then, we use another prompt-engineered transformer model, FeatureGPT (depicted in Figure 1(b)), to output a predicted probability (from 0 to 1) for each predefined feature category for a given video. Next, the similarity score, $Sim(V_{real},V_{sim})$, is calculated as a vector of differences between the predicted probabilities across the predefined categories:
@@
Although a single SCENIC script can produce multiple simulation videos with consistent core driving behaviors but variations in other parameters (e.g., road types or vehicle configurations), we utilize only a single video for similarity checks and iterative refinement, because this is sufficient to ensure that the generated scenario aligns with the essential driving behaviors of the original video. This approach provides the necessary focus for refining critical features while maintaining flexibility for generating diverse scenarios, enabling search-based testing.
@@
Experiments & Analysis
System Setup
@@
To construct the ScriptGPT model, we begin by writing 20 SCENIC scripts that cover diverse driving scenarios, such as overtaking, cruising, sudden stops due to obstacles, and turns in varying road and weather conditions. Using the SCENIC library and CARLA, we generate corresponding videos for each scenario and pair them with their respective SCENIC scripts, forming 20 $(V_{i},S_{i})$ pairs. These pairs are then used as input data for prompt engineering GPT-4o, selected for its ability to learn and generalize from a wide range of examples.
The empirical design choice we made is to design only 20 pairs because 20 scenarios are sufficient to cover most of the common interesting scenarios. Moreover, in practice, since GPT-4o API does not natively accept video formats like .mp4, we preprocess the videos by sampling frames and concatenating them into an n-dimensional array. The same preprocessing technique is applied during testing phases to ensure consistency between training and testing phases, and we also apply the same to the FeatureGPT.
@@
Preliminary Qualitative Result
@@
After completing the prompt engineering process, our framework is capable of automatically generating the 5 scenarios mentioned in Section Case Study without any human intervention during the testing phase. We also extended our framework to evaluate 50 randomly selected accidents from the CCD dataset, and found that 32 out of 50 (64%) scenarios generation can be fully automated without any human involvement, while 18 scenarios (36%) encountered syntax errors that required human debugging.
@@
Our framework significantly reduces the time required for real-to-simulation scenario generation. For the 32 videos that can be automatically generated, it takes 1.5 minutes per scenario on average during the testing phase, including iterative refinement, to produce a SCENIC script of approximately 70 lines of code. In contrast, manually coding and debugging a similar real-to-simulation scenario could take an experienced engineer several hours.
@@
The 5 scenarios in Section Case Study underwent 1-2 iterations of refinement, resulting in notable improvements in both accuracy and realism. In the extended evaluation of 50 accidents, iterative refinement happened in 17 scenarios (34%), further showcasing its potential to enhance scenario quality and be generalized to more crash scenarios.
@@
Limitations & Future Work
@@
Conclusion
In this paper, we have presented a novel framework for automatically converting real-world vehicle crash videos into simulation scenarios using prompt-engineered Video-Language Models. We have deploy multiple techniques including the similarity score metric and the iterative refinement process to ensure the generated scenarios closely align with the original videos. Despite the framework’s current limitations, such as reliance on data diversity and the challenges of handling complex or unseen scenarios, through multiple examples, it demonstrates clear potential for improving ADS testing. Future work will focus on expanding data diversity, conducting quantitative human studies, and extending the framework to other robotics domains.

### v2

@@ 1 Introduction
In this paper, we propose a novel framework that automates the conversion of real-world driving videos into detailed simulation scenarios. Our approach leverages prompt-engineered Video Language Models(VLMs) to transform dashcam videos into SCENIC scripts, enabling the automatic generation of realistic simulations in CARLA. In addition, rather than aiming solely to create an exact digital copy of the dashcam video, our framework prioritizes capturing the most essential driving behaviors while maintaining flexibility in parameters such as weather and road conditions. This flexibility supports search-based testing, a methodology that systematically varies parameters to identify edge cases and vulnerabilities in ADS. Furthermore, we introduce a similarity metric that iteratively refines the generated simulations by comparing key driving features between the real and simulated videos. The contributions of this work are fourfold: (1) an automated video-to-simulation pipeline that removes the need for manual scenario construction, (2) a similarity metric that bridge the gap between real and simulated scenarios, (3) an iterative feedback loop for scenario refinement using neural network-based feedback from VLM to capture core driving behaviors, and (4) a significant improvement in time efficiency, reducing scenario generation from hours to minutes while maintaining high fidelity to the original events.
2 Related Work
@@
In contrast, our framework integrates real-world driving data, such as dashcam crash videos, to generate simulation scenarios that not only capture critical driving behaviors but also provide flexibility for environmental parameters like road type and weather. Through few-shot prompt engineering, we tailor VLMs to create SCENIC scripts for scenario generation, bridging the gap between exact scenario replication and the need for diversity in simulation environments. This approach ensures realistic and behaviorally accurate test cases while supporting search-based testing for robust ADS validation.
3 Real-to-Sim Scenario Generation Framework
Given a real-life vehicle crash video (e.g., a dash camera recording), our objective is to generate corresponding simulation scenarios that accurately capture the core driving behaviors. In this paper, we define scenario as a driving event that includes environmental conditions (e.g., weather, road types), vehicle dynamics (e.g., speed, acceleration), and driving behaviors between entities (e.g., overtaking behavior or collisions). This approach enables robust and safe testing of Automated Driving Systems (ADS) by leveraging real-life close-call situations within a controlled simulation environment. Our framework consists of 4 components: (1) conversion of real-world video into SCENIC scripts, (2) generation of simulation videos from SCENIC scripts, (3) similarity analysis between the real and simulated videos, and (4) iterative refinement to ensure the simulated video’s consistency with the original scenario.
3.1 Video-to-Text Generation
@@
First, we convert the input video into a descriptive script called SCENIC. SCENIC is a probabilistic programming language designed for describing driving scenarios with a high level of precision, enabling detailed specification of environmental factors, road layouts, and vehicle behaviors for use in simulation platforms like CARLA. This conversion is accomplished using a prompt-engineered version of the pre-trained GPT-4o model.
@@
Although the initial prompt engineering process was conducted using simulation videos paired with their SCENIC scripts, through empirical testing, we found this approach could be extendable to real-world driving videos because the underlying visual and descriptive patterns (e.g., road layouts, traffic behaviors, and environmental factors) are consistent across both domains, allowing the model to effectively generalize its learned capabilities and accurately capture real-world scenarios.
3.2 Text-to-Video Generation
@@
3.3 Similarity Check
Next, we perform a similarity check on the simulated scenario $V_{sim}$ and the original $V_{real}$, measured on a set of predefined features. Ideally, the simulated and original scenario should closely match, allowing us to seamlessly replace the ego vehicle with any ADS (e.g. Baidu’s Apollo planner (Fan et al. 2018) and controller) for testing. However, due to the complexity of real-world scenarios, even with prompt engineering, discrepancies often arise, where the generated simulation may miss key features or introduce extra ones.
To address this, we introduce a similarity metric to ensure that the generated video captures the most important features from the original video. We predefine a set of crucial feature categories, such as the most critical and frequently encountered driving behaviors and environmental conditions, to examine the original crash video $V_{real}$ with the generated simulation $V_{sim}$. Then, we use another prompt-engineered transformer model, FeatureGPT (depicted in Figure 1(b)), to output a predicted probability (from 0 to 1) for each predefined feature category for a given video. Next, the similarity score vector, $Sim(V_{real},V_{sim})$, is calculated as a vector of differences between the predicted probabilities across the predefined categories:
@@ 3.4 Iterative Refinement
Although a single SCENIC script can produce multiple simulation videos with consistent core driving behaviors but variations in other parameters (e.g., road types or vehicle configurations), we utilize only a single video for similarity checks and iterative refinement, because multiple simulation videos has the same driving behaviors and refining on a video is sufficient to ensure that the generated scenario aligns with the essential driving behaviors of the original video. This approach provides the necessary focus for refining critical features while maintaining flexibility for generating diverse scenarios, enabling search-based testing.
@@
4 Experiments & Analysis
4.1 System Setup
@@
To construct the ScriptGPT model, we begin by carefully and manually writing 20 SCENIC scripts that cover diverse driving scenarios, such as overtaking, cruising, sudden stops due to obstacles, and turns in varying road and weather conditions. Using the SCENIC library and CARLA, we generate corresponding videos for each scenario and pair them with their respective SCENIC scripts, forming 20 $(V_{i},S_{i})$ pairs. These pairs are then used as input data for prompt engineering GPT-4o, selected for its ability to learn and generalize from a wide range of examples.
The empirical design choice we made is to design only 20 pairs because through empirical testing, 20 scenarios are sufficient to cover most of the common interesting scenarios. We believe we only need a very small number of prompt engineering due to the powerful generalization ability of pretrained GPT-4o. Moreover, in practice, since GPT-4o API does not natively accept video formats like .mp4, we preprocess the videos by sampling frames and concatenating them into an n-dimensional array. The same preprocessing technique is applied during testing phases to ensure consistency between training and testing phases, and we also apply the same to the FeatureGPT.
@@ 4.2 Case Study
4.3 Preliminary Qualitative Result
@@
After completing the prompt engineering process, our framework is capable of automatically generating the 5 scenarios mentioned in Section 4.2 without any human intervention during the testing phase. We also extended our framework to evaluate 50 randomly selected accidents from the CCD dataset, and found that 32 out of 50 (64%) scenarios generation can be fully automated without any human involvement, while 18 scenarios (36%) encountered errors (i.e. generating a keyword or object that is not supported by SCENIC syntax), thus cannot ran automatically.
@@
Our framework significantly reduces the time required for real-to-simulation scenario generation. For the 32 videos that can be automatically generated, it takes 1.5 minutes per scenario on average during the testing phase, including iterative refinement, to produce a SCENIC script of approximately 70 lines of code. In contrast, manually coding and debugging a similar real-to-simulation scenario could take an experienced engineer several hours. Even for the 36% scenarios that cannot automatically finish, it only takes minutes for human to fix the syntax error to make the pipeline work again.
@@
The 5 scenarios in Section 4.2 underwent 1-2 iterations of refinement, resulting in notable improvements in both accuracy and realism. In the extended evaluation of 50 accidents, iterative refinement happened in 17 scenarios (34%), further showcasing its potential to enhance scenario quality and be generalized to more crash scenarios.
@@
5 Limitations & Future Work
@@
6 Conclusion
In this paper, we have presented a novel framework for automatically converting real-world vehicle crash videos into simulation scenarios using prompt-engineered Video-Language Models. We have deploy multiple techniques including the similarity score vector metric and the iterative refinement process to ensure the generated scenarios closely align with the original videos. Despite the framework’s current limitations, such as reliance on data diversity and the challenges of handling complex or unseen scenarios, through multiple examples, it demonstrates clear potential for improving ADS testing. Future work will focus on expanding data diversity, conducting quantitative human studies, and extending the framework to other robotics domains.